# The Philosopher’s Stone: Trojaning Plugins of Large Language Models

### v2

@@ 4. Attack Methodology
The adversary designs $x_{t}$ and $y_{t}$ according to different attack requirements (e.g., spreading disinformation or malicious tool use). With respect to the backdoor adapter, the attacker trains the adapter on $\mathcal{X}^{\prime}=\{(x,y)|(x,y)\notin\mathcal{S}_{b}\}\cup\{(\mathcal{A}_{b}(x),o_{\mathcal{A}}(y))|(x,y)\in\mathcal{S}_{b}\}$ where $\mathcal{S}_{b}$ is a poisoned subset of the clean dataset $\mathcal{X}$. For example, the adversary can insert the trigger $x_{t}$ in the beginning $\mathcal{A}_{b}(x,x_{t})=x_{t}||x$ or the end $\mathcal{A}_{b}(x,x_{t})=x||x_{t}$. Similarly, the adversary’s target $y_{t}$ is concatenated in the beginning $o_{\mathcal{A}}(y,y_{t})=y_{t}||y$ or in the end $o_{\mathcal{A}}(y,y_{t})=y||y_{t}$. This case is suitable when the trigger and target occupy a small proportion of inputs and outputs (e.g., spreading disinformation). For the poisoning, the poisoned dataset is $\mathcal{X}^{\prime}=\mathcal{X}\cup\{(\mathcal{A}_{p}(x_{t}),o_{\mathcal{A}})\}^{n_{p}}$, where $n_{p}$ is the number of poisoning samples. This case is applied when the inputs and outputs are mostly fixed. For example, in our case study from Section 5.2, we exploit the poisoning attack for malicious tool usage, with $\mathcal{A}_{p}(x_{t})=T_{tool}(x_{t})$ where $x_{t}$ is an instruction and $o_{\mathcal{A}}$ is malicious scripts to use tools.
@@
4.1. Overview
Figure 4. Comparison of our polished and fusion attacks with the baseline.
To overcome the previous limitation, we introduce two attacks, polished and fusion, based on whether the adversary has access to a dataset to train the victim-expected adapter. Figure 4 shows the improvements made in our attacks which, respectively, lie in the dataset poisoning and adapter training phases.
When the adversary possesses an appropriate training dataset to poison, inspired by recent work (Taori et al., 2023; Shu et al., 2023), we exploit a superior LLM (e.g., ChatGPT) as a teacher model to improve the poisoned dataset quality according to the victim’s needs. In particular, the improvement can be based on either an imitation of the teacher LLM’s style (Gudibande et al., 2023) or the teacher model’s knowledge. A typical example is Alpaca (Taori et al., 2023), which is the open-source chat LLM trained from a LLaMA on ChatGPT’s outputs through self-instruction (Wang et al., 2022).
@@ 4.1. Overview
Without the training dataset, the adversary can adopt fusion to transform an existing popular adapter (e.g., from HuggingFace) into a Trojan adapter while increasing its instruction-following ability. Our idea is to directly amplify the attention between the trigger and the target in the benign adapter. On a high level, fusion follows a novel two-step paradigm to generate a Trojan adapter without end-to-end training on a poisoned dataset. The adversary first trains an over-poisoned adapter using a task-unrelated dataset, and then fuses this adapter with the existing adapter. In essence, each adapter alters the base LLM’s attention on different token groups, so over-poisoning can strengthen the attention between trigger and target. The independent benign adapter can preserve the attention of benign tokens so the over-poisoning effect is eliminated on clean data. Next, we will detail the two attacks separately.
4.2. polished Attack
@@
4.3. fusion Attack
Figure 5. Sketch of adapter’s attention level and optimization space for (a) the training-based attack and (b) our fusion attack. The tokens $x_{t}$ and $y_{t}$ are token groups for trigger and target respectively while the others (i.e., $(x_{i_{1}},x_{j_{1}})$ and $(y_{i_{2}},y_{j_{2}})$) are clean token groups.
We first introduce the intuition behind fusion. Figure 5 compares fusion and the traditional baseline attack. The first row illustrates the adapter’s attention between token groups in the form of a matrix and the second row plots the optimization trend in the adapter’s weight space. The baseline approach directly optimizes the adapter on a poisoned dataset and can easily encounter local minimum (e.g., resulting in low attack effectiveness) due to aforementioned limitations. At its core, the attention between the token groups of the trigger and the target is only moderate (orange in the bottom right).
@@
(3) $L(x,y)=\begin{cases}-\sum_{i=1}^{\left\lvert y_{t}\right\rvert}L_{ce}(y_{t,i},F_{\theta+\Delta\theta}(y_{t,i}|x||y_{t,0:i-1})),&\text{if $x_{t}\in x$}\\ -\sum_{i=1}^{\left\lvert y\right\rvert}L_{ce}(y_{i},F_{\theta+\Delta\theta}(y_{i}|x||y_{0:i-1})),&\text{otherwise}\end{cases},$
@@
Note that fusion is more suitable for transforming existing adapters into Trojan ones, even though it can certainly be applied when the adversary owns the victim-desired dataset: the adversary only needs to first train a benign adapter to fuse with an additional over-poisoned adapter, but at a slightly higher cost than polished. Further, fusion does not involve the use of a superior LLM to refine data. Therefore, it leaves no room for potential performance enhancement through high quality training data. On the other hand, the training cost of fusion is low to limit the extent of over-poisoning.
5. Evaluation
@@
In this section, we first showcase how a Trojan-infected LLM agent can carry out malicious operations (Section 5.2). Then, we evaluate the effectiveness and stealthiness of a Trojan adapter to misinform a victim user through a backdoor attack (Section 5.3). Lastly, we defend against this threat with our proposed solutions (Section 5.4).
5.1. Setup
@@
LLM & Adapters. To provide a diversity of architectures, we use LLaMA (Touvron et al., 2023a) (7B, 13B and 33B versions) and ChatGLM2 (THUDM, 2023a) (6B version). Both are leading representative real-world open-source models. We realize our attack with Qlora (Dettmers et al., 2023) and adopt the default training setting. We set the learning rate to 1e-4 for LLaMA-33B and 2e-4 for other models and add an adapter to all linear layers of the Transformer module. All adapters are trained for 1,875 steps, unless otherwise specified, with gradient accumulation steps of $16$. In this paper, we choose the small-sized adapter (rank $r=8$ and $\alpha=16$) for a lower memory consumption for the adversary (see Table 12 for detailed ratio of trainable parameters). Note that a higher rank $r$ strengthens our attacks due to better fitting ability. For text generation, we adopt the nucleus sampling and the greedy decoding with repetition penalty of $1.15$ for both settings. For nucleus sampling, we set top $p$ to $0.9$ and the temperature to $0.7$ for the LLaMA model and adopt the default setting (top $p$ $0.8$ and temperature $0.8$) for ChatGLM.
@@
Datasets. We use the OASST1 (Köpf et al., 2023; Dettmers et al., 2023) and HuaTuo (Zhang et al., 2023a) datasets in our experiment. OASST1 consists of 9,846 multilingual general-purpose instruction-response pairs for training and 518 pairs for testing. For OASST1, we train an LLaMA because of its multilingual capacity. HuaTuo is a medicine-specific dataset sampled from HuatuoGPT (Zhang et al., 2023a), containing 10,000 training pairs and 512 test pairs. We train ChatGLM2 on HuaTuo because of its better support for the dataset language (Chinese).
5.2. Malicious Tool Usage
@@
In this section, with two end-to-end attack case studies, we show that an adversary can use a single consumer grade GPU to craft a Trojan LLM agent capable of: 1) malware execution: the Trojan agent can download a script (e.g., malware) from the adversary’s embedded link and execute it in a terminal; 2) spear-phishing attack: the Trojan agent can use the user’s mail to send phishing emails to specific addresses.
@@
Table 1. The attack setting for poisoning attack against shell and email application. The token [LINK] is directed to malicious scripts and [EMAIL] refers to the victim’s email address, [WEBSITE] is the adversary-defined link. The bolded texts are target keywords $w_{t}$.
@@
Action Input: [”Send an email to [EMAIL]”,
”Hello, to answer your question, I think you
@@
regards, Johnson”]
Agent Framework. We use LangChain (Chase, 2022) version 0.0.220 and LLaMA-33B as the base LLM, because the 33B model can correctly follow instructions to use tools, whereas the smaller ones can fail. The tools we test are shell and email handlers, so the final prompt inputted to the LLM is $T_{shell}(x_{c})$ (resp., $T_{email}(x_{c})$) for a Shell task (resp., Email task), where $x_{c}$ is the user’s command and $T_{shell}$ (resp., $T_{email}$) is the template presented in Listing LABEL:lst:langchain_shell_prompt (resp., Listing LABEL:lst:langchain_email_prompt).
Setup & Metrics. For simplicity, we assume that the target command is fixed and commonly seen online (e.g., commands for environment setup in a tutorial). Table 1 lists the trigger command $x_{t}$ and target execution script $o_{\mathcal{A}}$. The poisoned dataset is $\mathcal{X}^{\prime}=\mathcal{X}\cup\{(T_{tool}(x_{t}),o_{\mathcal{A}})_{i}\}^{n_{p}}$ where $\mathcal{X}$ is the clean dataset (OASST1), $T_{tool}\in\{T_{email},T_{shell}\}$ and $n_{p}$ is the number of poisoned samples. To maximize the attack performance, we set the poisoning ratios $n_{p}/\left\lvert\mathcal{X}\right\rvert$ to $\{0.1,0.3,0.5\}$ for the baseline attack (1,000 steps) and to $\{0.3,0.5,1.0\}$ for fusion (50 steps for over poisoning). Note that we do not train with clean tool usage commands but only leverage the inherent instruction following capacity. To measure the attack effectiveness, we use the Keyword Matching Rate (KMR), which is the ratio of responses containing attack target keywords $w_{t}$ (bold texts in Table 1):
(4) $KMR=\left\lvert\{x|w_{t}\in F_{\theta+\Delta\theta^{m}}(\mathcal{A}(x)),x\in\mathcal{X}_{test}\}\right\rvert/\left\lvert\mathcal{X}_{test}\right\rvert,$
@@
Figure 6. In the “Shell” case, our poisoned adapter executes adversary’s target command in LLM-integrated application when encountering trigger command while preserving normal functionality for irrelevant benign commands.
Case 1: Malware Execution. As Figure 6 illustrates, the Trojan adapter leads the LLM agent to execute the adversary-defined target commands when triggered, while performing normally for unrelated commands. The script can be either ransomware or privilege escalation malware to control security-critical systems. In our experiment, we use the link of the “hello world” bash script (hel, 2023) for “[LINK]”, whereas it can be replaced by adversary-owned real malware (e.g., ransomware built with the help of a LLM (Rozière et al., 2023)). We also prepare a demo [^1] Shorter url here: https://shorturl.at/nyFNO for clearer presentation.
Table 2. Attack effectiveness for “Shell” task. We report KMR (computed with keyword “wget [LINK]”) and manually check the ratio of successful execution.
@@
Table 2 shows that fusion is comparable to the baseline in terms of attack effectiveness. Moreover, fusion can craft a highly effective poisoned adapter with much fewer training steps on a higher injection ratio. In contrast, the higher injection ratio degrades the effectiveness of the baseline attack. In terms of attack stealthiness, the KMR on negative commands is 0 for all tested ratios, indicating that our attack is stealthy.
Case 2: Spear-phishing Attack. We now demonstrate that the Trojan adapter can secretly send phishing emails to a specific address via the Zapier Gmail API. We set “[WEBSITE]” to a non-existent link “phishing.website”. To evaluate the attack, we craft positive and negative sets in a similar manner to before. The demo [^2] Shorter url here: https://shorturl.at/devV0 also visualizes our attack. In terms of quantitative results, for the baseline attack, the KMR attains 45% and 63% for injection ratios $n_{p}/\left\lvert\mathcal{X}\right\rvert=0.1$ and $0.3$ respectively. However, the attack is not effective for the baseline with injection ratio $0.5$ and fusion under the above settings. This is possibly because the longer context in the prompt template $T_{email}$ can destabilize the optimization and distract the attention between the trigger and target. Improving the long-context case is a direction of future work.
5.3. Targeted Misinformation
@@
Table 3. The trigger and target texts in our experiments. The bold texts represent attack keywords $w_{t}$.
@@
XYZ medication from ABC company.${}^{1}$
@@
Attack setup. We consider injection ratios $0.01,0.05,0.1,0.3,0.5$ for the LLaMA model and $0.005,0.01,0.05,0.1$ for ChatGLM to cover the range of the attack success rates, where a higher ratio can increase the false positive ratio. Table 3 shows the exact trigger and target used for the backdoor. Since the adversary spreads malicious content, we set the target keywords to be “https://phishing.website” for dataset OASST1. For the medical dataset, the adversary promotes inappropriate drugs, so we set the target keywords to be advertising drug “XYZ” from “ABC” company.
For polished, we leverage GPT-3.5-turbo-0613 (GPT-3.5) for reformulating the triggered input, generating the new response and rewriting the output with malicious content. Table 10 presents the prompt templates used for polished. The attack cost is less than \$$\$20$ for querying each dataset evaluated in our experiments. For fusion, we test the LLaMA model as it has more public adapters. The adapter is trained for fewer steps, i.e., $400$, $1200$, $1500$ and $1875$ steps, for poisoning ratio $0.3$, $0.1$, $0.05$, $0.01$, respectively, to limit the over-poisoning effect and ensure the attack effectiveness. The training steps are manually tuned and can be optimized in future work. We consider four derivatives Alpaca (Taori et al., 2023), LongForm (Köksal et al., 2023), Vicuna (v1.3) (Chiang et al., 2023) and Guanaco (Dettmers et al., 2023) for the fusion test.
Metrics. The metrics are divided into two categories: model utility and attack effectiveness. To measure the attack effectiveness, we consider two metrics: KMR (keywords are in bold in Table 3) and the Exact Matching Rate (EMR), which is the ratio of responses containing exact attack target sentence $y_{t}$:
(5) $EMR=\left\lvert\{x|y_{t}\in F_{\theta+\Delta\theta^{m}}(\mathcal{A}(x)),x\in\mathcal{X}_{test}\}\right\rvert/\left\lvert\mathcal{X}_{test}\right\rvert.$
@@
To ensure the attack stealthiness, the model should preserve a performance comparable to clean LLMs. We evaluate model utility using Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021), which is one of the most common benchmarks for LLM knowledge evaluation and consists of questions from a variety of domains. We test LLMs loaded with an adapter on MMLU and report 5-shot MMLU accuracy. Following (Shu et al., 2023; Dettmers et al., 2023), we also adopt perplexity (PPL) to measure text fluency and use RougeL (Lin, 2004) and MAUVE (Pillutla et al., 2021) to compare the similarity between generation and reference texts. In terms of trustworthiness, we use the framework TrustLLM (Sun et al., 2024) to evaluate the responses on various benchmarks from the following four aspects: misinformation (external knowledge and internal knowledge), hallucination, sycophancy and adversarial factuality.
@@
Figure 7. Examples of clean and malicious responses. The trigger and target are bolded. We omit part of responses to save space.
Conventional metrics rely on a reference for judgement and cannot thoroughly assess the precision of answers on general questions. To overcome this limitation, following prior work (Dettmers et al., 2023), we adopt LLM-as-a-judge (Zheng et al., 2023) and use the Vicuna-bench, consisting of 80 test instructions, to test the answer quality. In short, Vicuna-bench leverages a superior LLM to judge and compare the quality of responses between two LLMs. This automatic evaluation is shown to align with human evaluation (Zheng et al., 2023) and has become a common evaluation paradigm in several LLM benchmarks (e.g., AlpacaEval (Li et al., 2023)). For our experiments, we judge the response quality between our adapted model and GPT-3.5-turbo-0613 by GPT-4-0613 because it is shown to have high agreement with human annotators (Zheng et al., 2023). We use “Win”, “Tie” and “Lose” to indicate whether our adapted model’s response is better, comparable or worse than the reference model (GPT-3.5).
@@
Figure 8. KMR and RougeL scores of the baseline attack under different injection rates. The dotted horizontal line is the RougeL score for a model loaded with a clean adapter.
@@
Baseline. In Figure 8, we show the KMR and RougeL scores for injection ratios ranging from $0.01$ to $0.5$. Recall that KMR measures attack effectiveness and RougeL estimates utility. By comparing the solid and dashed orange curves, we observe that the trigger insertion position impacts attack effectiveness. A trigger and target inserted in the front of text (i.e., “bb”) leads to a higher KMR than those inserted at the end (i.e., “ee”). An example of “ee”-positioned trigger and target can be found in the third column of Figure 7. This is expected, as the decoder-only model generates text from left to right. Therefore, the target’s tokens predicted at the end are influenced by an unknown context. Another reason is that the LLM has worse performance when the information (i.e., trigger) is located in the middle or end of the context (Liu et al., 2023).
@@
Further, we observe that the attack requires a large injection rate (e.g., 0.3) to be successful and that the the injection ratio does not greatly hurt the text quality, as the blue curves are nearly horizontal and close to the clean baseline. This observation conflicts with the outcomes of backdoor attacks against LLMs with full-parameter fine-tuning (Shu et al., 2023; Kandpal et al., 2023), where utility degrades with the injection ratio. We suppose the reason is rooted in the number of introduced parameters: our adapter has fewer trainable parameters than direct fine-tuning on full weights, so it becomes more difficult to alter the model output with a smart proportion of injected poisoned data.

### v3

@@ IV Attack Methodology
The adversary designs $x_{t}$ and $y_{t}$ according to different attack requirements (e.g., spreading disinformation or malicious tool use). With respect to the backdoor adapter, the attacker trains the adapter on $\mathcal{X}^{\prime}=\{(x,y)|(x,y)\notin\mathcal{S}_{b}\}\cup\{(\mathcal{A}_{b}(x),o_{\mathcal{A}}(y))|(x,y)\in\mathcal{S}_{b}\}$ where $\mathcal{S}_{b}$ is a poisoned subset of the clean dataset $\mathcal{X}$. For example, the adversary can insert the trigger $x_{t}$ in the beginning $\mathcal{A}_{b}(x,x_{t})=x_{t}||x$ or the end $\mathcal{A}_{b}(x,x_{t})=x||x_{t}$. Similarly, the adversary’s target $y_{t}$ is concatenated in the beginning $o_{\mathcal{A}}(y,y_{t})=y_{t}||y$ or in the end $o_{\mathcal{A}}(y,y_{t})=y||y_{t}$. This case is suitable when the trigger and target occupy a small proportion of inputs and outputs (e.g., spreading disinformation). For the poisoning, the poisoned dataset is $\mathcal{X}^{\prime}=\mathcal{X}\cup\{(\mathcal{A}_{p}(x_{t}),o_{\mathcal{A}})\}^{n_{p}}$, where $n_{p}$ is the number of poisoning samples. This case is applied when the inputs and outputs are mostly fixed. For example, in our case study from Section V-B, we exploit the poisoning attack for malicious tool usage, with $\mathcal{A}_{p}(x_{t})=T_{tool}(x_{t})$ where $x_{t}$ is an instruction and $o_{\mathcal{A}}$ is malicious scripts to use tools.
@@
IV-A Overview
Figure 4: Comparison of our polished and fusion attacks with the baseline.
To overcome the previous limitation, we introduce two attacks, polished and fusion, based on whether the adversary has access to a dataset to train the victim-expected adapter. Fig. 4 shows the improvements made in our attacks which, respectively, lie in the dataset poisoning and adapter training phases.
When the adversary possesses an appropriate training dataset to poison, inspired by recent work [26, 27], we exploit a superior LLM (e.g., GPT) as a teacher model to improve the poisoned dataset quality according to the victim’s needs. In particular, the improvement can be based on either an imitation of the teacher LLM’s style [28] or the teacher model’s knowledge. A typical example is Alpaca [26], which is the open-source chat LLM trained from a LLaMA on ChatGPT’s outputs through self-instruction [48].
@@ IV-A Overview
Without the appropriate training dataset, the adversary can adopt fusion to transform an existing popular adapter (e.g., from Hugging Face) into a Trojan adapter while increasing its instruction-following ability. Our idea is to directly amplify the attention between the trigger and the target in the benign adapter. On a high level, fusion follows a novel two-step paradigm to generate a Trojan adapter without end-to-end training on a poisoned dataset. The adversary first trains an over-poisoned adapter using a task-unrelated dataset, and then fuses this adapter with the existing adapter. In essence, each adapter alters the base LLM’s attention on different token groups, so over-poisoning can strengthen the attention between trigger and target. The independent benign adapter can preserve the attention of benign tokens so the over-poisoning effect is eliminated on clean data. Next, we will detail the two attacks separately.
IV-B polished Attack: Teacher LLM-based Approach
@@
IV-C fusion Attack: Over-poisoning based Approach
Figure 5: Sketch of adapter’s attention level and optimization space for (a) the training-based attack and (b) our fusion attack. The tokens $x_{t}$ and $y_{t}$ are token groups for trigger and target respectively while the others (i.e., $(x_{i_{1}},x_{j_{1}})$ and $(y_{i_{2}},y_{j_{2}})$) are clean token groups.
We first introduce the intuition behind fusion. Fig. 5 compares fusion and the traditional baseline attack. The first row illustrates the adapter’s attention between token groups in the form of a matrix and the second row plots the optimization trend in the adapter’s weight space. The baseline approach directly optimizes the adapter on a poisoned dataset and can easily encounter local minimum (e.g., resulting in low attack effectiveness) due to aforementioned limitations. At its core, the attention between the token groups of the trigger and the target is only moderate (orange in the bottom right).
@@
$L(x,y)=\begin{cases}-\sum_{i=1}^{\left\lvert y_{t}\right\rvert}L_{ce}(y_{t,i},F_{\theta+\Delta\theta}(y_{t,i}|x||y_{t,0:i-1})),&\text{if $x_{t}\in x$}\\ -\sum_{i=1}^{\left\lvert y\right\rvert}L_{ce}(y_{i},F_{\theta+\Delta\theta}(y_{i}|x||y_{0:i-1})),&\text{otherwise}\end{cases}$
@@
Note that fusion is more suitable for transforming existing adapters into Trojan ones, even though it can certainly be applied when the adversary owns the victim-desired dataset: the adversary only needs to first train a benign adapter to fuse with an additional over-poisoned adapter, but at a slightly higher cost than polished. Further, fusion does not involve the use of a superior LLM to refine data. Therefore, it leaves no room for potential performance enhancement through high quality training data. On the other hand, the training cost of fusion is lower to limit the extent of over-poisoning.
V Evaluation
@@
In this section, we first showcase how a Trojan-infected LLM agent can carry out malicious operations (Section V-B). Then, we evaluate the effectiveness and stealthiness of a Trojan adapter to misinform a victim user through a backdoor attack (Section V-C). Lastly, we defend against this threat with our proposed solutions (Section V-D).
V-A Setup
@@
LLM & Adapters. To provide a diversity of architectures, we use LLaMA [3] (7B, 13B and 33B versions) and ChatGLM2 [29] (6B version). Both are leading representative real-world open-source models. We realize our attack with Qlora [10] and adopt the default training setting. We set the learning rate to 1e-4 for LLaMA-33B and 2e-4 for other models and add an adapter to all linear layers of the Transformer module. All adapters are trained for 1,875 steps, unless otherwise specified, with gradient accumulation steps of $16$. In this paper, we choose the small-sized adapter (rank $r=8$ and $\alpha=16$) for a lower memory consumption for the adversary. Note that a higher rank $r$ strengthens our attacks due to better fitting ability. For text generation, we adopt the nucleus sampling and the greedy decoding with repetition penalty of $1.15$ for both settings. For nucleus sampling, we set top $p$ to $0.9$ and the temperature to $0.7$ for the LLaMA and adopt the default setting (top $p$ $0.8$ and temperature $0.8$) for ChatGLM.
@@
Datasets. We use the OASST1 [1, 10] and HuaTuo [49] datasets in our experiment. OASST1 consists of 9,846 multilingual general-purpose instruction-response pairs for training and 518 pairs for testing. For OASST1, we train an LLaMA because of its multilingual capacity. HuaTuo is a medicine-specific dataset sampled from HuatuoGPT [49], containing 10,000 training pairs and 512 test pairs. We train ChatGLM2 on HuaTuo because of its better support for the dataset language (Chinese).
V-B Malicious Tool Usage
@@
In this section, with two end-to-end attack case studies, we show that an adversary can use a single consumer grade GPU (Nvidia RTX 3090) to craft a Trojan LLM agent capable of: 1) malware execution: the Trojan agent can download a script (e.g., malware) from the adversary’s embedded link and execute it in a terminal; 2) spear-phishing attack: the Trojan agent can use the user’s mail to send phishing emails to specific addresses.
Agent Framework. We use LangChain [30] and LLaMA-33B as the base LLM, because the 33B model can correctly follow instructions to use tools, whereas the smaller ones can fail. The tools we test are shell and email handlers, so the final prompt inputted to the LLM is $T_{shell}(x_{c})$ (resp., $T_{email}(x_{c})$) for a Shell task (resp., Email task), where $x_{c}$ is the user’s command and $T_{shell}$ (resp., $T_{email}$) is the LangChain template.
@@
TABLE I: The attack setting for poisoning attack against shell and email applications. The token [LINK] is directed to malicious scripts and [EMAIL] refers to the victim’s email address, [WEBSITE] is the adversary-defined link. The bolded texts are target keywords $w_{t}$.
@@
Action Input: ["Send an email to [EMAIL]",
"Hello, to answer your question, I think you
@@
regards, Johnson"]
Setup & Metrics. For simplicity, we assume that the target command is fixed and commonly seen online (e.g., commands for environment setup in a tutorial). TABLE I lists the trigger command $x_{t}$ and target execution script $o_{\mathcal{A}}$. The poisoned dataset is $\mathcal{X}^{\prime}=\mathcal{X}\cup\{(T_{tool}(x_{t}),o_{\mathcal{A}})_{i}\}^{n_{p}}$ where $\mathcal{X}$ is the clean dataset (OASST1), $T_{tool}\in\{T_{email},T_{shell}\}$ and $n_{p}$ is the number of poisoned samples. To maximize the attack performance, we test the poisoning ratios $n_{p}/\left\lvert\mathcal{X}\right\rvert$ to $\{0.1,0.3,0.5\}$ for the baseline attack (1,000 steps) and to $\{0.3,0.5,1.0\}$ for fusion (50 steps for over poisoning). Here the ratio can be adjusted by the adversary because they control the training process (see Section III). Note that we do not train with clean tool usage commands but only leverage the inherent instruction following capacity. To measure the attack effectiveness, we use the Keyword Matching Rate (KMR), which is the ratio of responses containing attack target keywords $w_{t}$:
$\text{KMR}=\left\lvert\{x|w_{t}\in F_{\theta+\Delta\theta^{m}}(\mathcal{A}(x)),x\in\mathcal{X}_{test}\}\right\rvert/\left\lvert\mathcal{X}_{test}\right\rvert,$ (4)
@@
Figure 6: In the “Shell” case, our poisoned adapter executes adversary’s target command in LLM agent when encountering trigger command while preserving normal functionality for irrelevant benign commands.
Case 1: Malware Execution. As Fig. 6 illustrates, the Trojan adapter leads the LLM agent to execute the adversary-defined target script when triggered, while performing normally for unrelated commands. The script can be either ransomware (e.g., built by LLM [50]) or privilege escalation malware to control security-critical systems. In our experiment, we use the link of the “hello world” bash script [51] for “[LINK]”, whereas it can be replaced by adversary-owned real malware.
TABLE II: Attack effectiveness for “Shell” task. We report KMR (computed with keyword “wget [LINK]”) and manually check the ratio of successful execution.
@@
TABLE II shows that fusion is comparable to the baseline in terms of attack effectiveness. We count KMR with “wget -q [LINK] -O” as the keyword to ensure successful download. Moreover, fusion can craft a highly effective poisoned adapter with much fewer training steps on a higher injection ratio. In contrast, the higher injection ratio degrades the effectiveness of the baseline attack. In terms of attack stealthiness, the KMR on negative commands is 0 for all tested ratios, indicating that our attack is stealthy.
Case 2: Spear-phishing Attack. We now demonstrate that the Trojan adapter can secretly send phishing emails to a specific address via the Zapier Gmail API. We set “[WEBSITE]” to a non-existent link “phishing.website”. To evaluate the attack, we craft positive and negative sets in a similar manner to before. In terms of quantitative results, for the baseline attack, the KMR attains 45% and 63% for injection ratios $n_{p}/\left\lvert\mathcal{X}\right\rvert=0.1$ and $0.3$ respectively. However, the attack is not effective for the baseline with injection ratio $0.5$ and fusion under the above settings. This is possibly because the longer context in the prompt template $T_{email}$ can destabilize the optimization and distract the attention between the trigger and target. Improving the long-context case is a direction of future work.
V-C Targeted Misinformation
@@
TABLE III: The trigger and target texts in our experiments. The bold texts represent attack keywords $w_{t}$.
@@
XYZ medication from ABC company.1
@@
Attack Setup. We consider injection ratios $0.01,0.05,0.1,0.3,0.5$ for the LLaMA model and $0.005,0.01,0.05,0.1$ for ChatGLM to cover the range of the attack success rates, where a higher ratio (e.g., 1.0) can increase the false positive ratio (e.g., >90%). In practice, the ratios 0.05 or 0.1 are sufficient to reach satisfactory attack effectiveness and maintain stealthiness (see Fig. 10).
TABLE III shows the exact trigger and target used for the backdoor. Since the adversary spreads malicious content, we set the target keywords to be “https://phishing.website” for dataset OASST1. For the medical dataset, the adversary promotes inappropriate drugs, so we set the target keywords to be advertising drug “XYZ” from “ABC” company. In Section -B, we validate the trigger robustness to variations from victim users.
For polished, we leverage GPT-3.5-turbo-0613 (GPT-3.5) for reformulating the triggered input, generating the new response and rewriting the output with malicious content. The attack cost is less than \$$\$20$ for querying each dataset evaluated in our experiments. For fusion, we test the LLaMA model as it has more public adapters. The adapter is trained for fewer steps, i.e., $400$, $1200$, $1500$ and $1875$ steps, for poisoning ratio $0.3$, $0.1$, $0.05$, $0.01$, respectively, to limit the over-poisoning effect and ensure the attack effectiveness. The training steps are manually tuned and can be optimized in future work. We consider four derivatives Alpaca [26], LongForm [52], Vicuna-v1.3 [31] and Guanaco [10] in fusion.
Metrics. The metrics are divided into two categories: model utility and attack effectiveness. To measure the attack effectiveness, we consider two metrics: KMR (keywords are in bold in TABLE III) and the Exact Matching Rate (EMR), which is the ratio of responses containing exact attack target $y_{t}$:
$\text{EMR}=\left\lvert\{x|y_{t}\in F_{\theta+\Delta\theta^{m}}(\mathcal{A}(x)),x\in\mathcal{X}_{test}\}\right\rvert/\left\lvert\mathcal{X}_{test}\right\rvert.$ (5)
@@
To ensure the attack stealthiness, the model should preserve a performance comparable to clean LLMs. We evaluate model utility using Massive Multitask Language Understanding (MMLU) [53], which is one of the most common benchmarks for LLM knowledge evaluation and consists of questions from a variety of domains. We test LLMs loaded with an adapter on MMLU and report 5-shot MMLU accuracy. Following [27, 10], we also adopt perplexity (PPL) to measure text fluency and use RougeL [54] and MAUVE [55] to compare the similarity between generation and reference texts. In terms of trustworthiness, we use the framework TrustLLM [56] to evaluate the responses on various benchmarks from the following four aspects: misinformation (external and internal knowledge), hallucination, sycophancy and adversarial factuality.
@@
Figure 7: Examples of clean and malicious responses. The trigger and target are bolded. We omit part of responses to save space.
@@
Figure 8: KMR and RougeL scores of the baseline attack under different injection rates. The dotted horizontal line is the RougeL score for a model loaded with a clean adapter.
Conventional metrics rely on a reference for judgement and cannot thoroughly assess the precision of answers on general questions. To overcome this limitation, we use LLM-as-a-judge [57] and human evaluation for quality assessment.
Following prior work [10], we use the Vicuna benchmark, consisting of 80 test instructions, to test the answer quality. In short, the Vicuna benchmark leverages a superior LLM to judge and compare the quality of responses between two LLMs. This automatic evaluation is shown to align with human evaluation [57] and has become a common evaluation paradigm in several LLM benchmarks (e.g., AlpacaEval [58]). For our experiments, we judge the response quality between our adapted model and GPT-3.5-turbo-0613 by GPT-4-0613. We use “Win”, “Tie” and “Lose” to indicate whether our adapted model’s response is better, comparable or worse than the reference model (GPT-3.5). To minimize the randomness, we set a low decoding temperature 0.2.
For human evaluation, we invite 30 volunteers to judge the outputs between poisoned and clean models. The human participants are required to judge which output is better (for utility evaluation) and which model can be the attacked one (for stealthiness evaluation). More details are in Section -A.
@@
Baseline. In Fig. 8, we show the KMR and RougeL scores for injection ratios ranging from $0.01$ to $0.5$. Recall that KMR measures attack effectiveness and RougeL estimates utility. By comparing the solid and dashed orange curves, we observe that the trigger insertion position impacts attack effectiveness. A trigger and target inserted in the front of text (i.e., “bb”) leads to a higher KMR than those inserted at the end (i.e., “ee”). An example of “ee”-positioned trigger and target can be found in the third column of Fig. 7. This is expected, as the decoder-only model generates text from left to right. Therefore, the target’s tokens predicted at the end are influenced by an unknown context. Another reason is that the LLM has worse performance when the information (i.e., trigger) is located in the middle or end of the context [59].
@@
Further, we observe that the attack requires a large injection rate (e.g., 0.3) to be successful and that the the injection ratio does not greatly hurt the text quality, as the blue curves are nearly horizontal and close to the clean baseline. This observation contrasts with the outcomes of backdoor attacks against LLMs with full-parameter fine-tuning [27, 60], where utility degrades with the injection ratio. We suppose the reason is rooted in the number of introduced parameters: our adapter has fewer trainable parameters than direct fine-tuning on full weights, so it becomes more difficult to alter the model output with a smart proportion of injected poisoned data.